### v1

# Towards Enforcing Company Policy Adherence in Agentic Workflows

Abstract
Large Language Model (LLM) agents hold promise for a flexible and scalable alternative to traditional business process automation, but struggle to reliably follow complex company policies. In this study we introduce a deterministic, transparent, and modular framework for enforcing business policy adherence in agentic workflows. Our method operates in two phases: (1) an offline buildtime stage that compiles policy documents into verifiable guard code associated with tool use, and (2) a runtime integration where these guards ensure compliance before each agent action. We demonstrate our approach on the challenging $\tau$-bench Airlines domain, showing encouraging preliminary results in policy enforcement, and further outline key challenges for real-world deployments.
Towards Enforcing Company Policy Adherence in Agentic Workflows
Naama Zwerdling          David Boaz          Ella Rabinovich          Guy Uziel David Amid                  Ateret Anaby-Tavor IBM Research naamaz@il.ibm.com, davidbo@il.ibm.com, ella.rabinovich1@ibm.com
1 Introduction
Large Language Models are reshaping artificial intelligence, shifting from static language processors to dynamic, task-oriented agents capable of planning, executing, and refining their actions. These agents hold the potential for transformative applications across various domains, including healthcare (Abbasian et al., 2023; Mehandru et al., 2024), finance (Li et al., 2024; Xiao et al., 2024; Ding et al., 2024), education Yang et al. (2024); Xu et al. (2024), and customer support Huang et al. (2024); Rome et al. (2024). LLM agents have been revolutionarily positioned as routing systems that act and perform tasks with minimal human intervention.
Agentic AI for business process automation refers to the increasing use of LLM-based agents to plan, make decisions, and carry out tasks within business workflows—often spanning multiple steps and following business constraints—without the need for constant human oversight. For example, modifying a flight reservation (such as changing a seat) may involve a sequence of actions: checking customer eligibility, retrieving available seats, and updating the reservation details if all conditions are met. Canceling a reservation with a refund might depend on the presence of insurance, while offering compensation for a company-initiated cancellation should occur only per an explicit customer request. These constraints are typically specified in a company’s policy documents and have traditionally been enforced by human service agents or hard-coded by domain experts into customer support systems such as IBM Watson Assistant, Microsoft Power Automate, or Google Dialogflow.
Agentic systems carry the promise to replace deterministic, carefuly-designed and hard-coded business processes, with efficient, more flexible and easily maintainable solutions. Recent studies have shown steady progress in the ability of agentic systems to complete complex, multi-step, and multi-turn tasks using a predefined set of tools (Patil et al., 2024; Zhou et al., 2023; Huang et al., 2024). However, agents perform poorly when required to follow policies involving sophisticated navigation or branching logic according to company guidelines (Yao et al., 2024; Li et al., 2025).
Despite the clear practical significance, current approaches to agentic policy adherence rely on the "best-effort" strategy: appending policy documents to the agent’s prompt and instructing it to navigate business flow while complying with the guidelines – an inherently non-deterministic methodology, that is susceptible to attacks and does not scale effectively (Yao et al., 2024; Li et al., 2025). We argue that the ability to reliably follow policies constitutes the ultimate test, a "make-or-break" factor in the adoption of agentic AI at the enterprise scale. We propose an automatic end-to-end solution for enforcement of company policy guidelines in agentic workflows; [^1] Policies addressed in this study are those directly protecting tool invocation; see Limitations section for details. solution that is deterministic, predictable, transparent and requires only limited domain expert intervention during its buildtime phase.
Figure 1: During the offline buildtime step (left), policy documents, along with system schema and tools specification are compiled into compact, easily interpretable mapping of concrete policies onto tools (editable json in this case); which in turn is used to generate ToolGuards – code that verifies that policies hold given a system snapshot. ToolGuards are integrated into agentic runtime at the tool invocation point (right). We mark two possible points of domain-expert intervention: reviewing textual Tool-Policy mapper outcome, and the generated ToolGuards code.
Our approach enforces policy adherence through a two-phase process illustrated in Figure 1: (1) an offline buildtime step automatically maps policy fragments to the relevant tools and generates policy validation code; (2) during runtime these validators—called ToolGuards—are executed before agent’s tool invocation. If a planned action violates a policy, the agent is prompted to self-reflect and revise its plan before proceeding. Ultimately, the deployed ToolGuards will prevent the agent from taking an action violating a policy.
We evaluate the proposed approach on the $\tau$-bench Airlines domain (Yao et al., 2024) with a large set of (naturalistic, real-world) policies, 14 tools, and diverse customer-service scenarios. [^2] The other $\tau$-bench domain of Retail is smaller and less diverse; hence we limit the scope of this work to Airlines. We show that this ambitious and challenging task pushes the limits of contemporary LLMs, present encouraging preliminary results, and outline directions for future research and practical steps needed for this approach to mature toward enterprise adoption. Our code and data will be made available.
2 Related Work
Prior art in the area of constraint and policy adherence in agentic workflows is very sparse, with two major studies, to the best of our knowledge, explicitly addressing the integration of company policies within agentic flows (contrary to post-hoc remedy). Yao et al. (2024) were among the pioneers in this field, introducing a framework for evaluation of company policies adherence in multi-turn simulated flows in two domains: Airlines and Retail. Policies are represented by short documents, where domain-specific rules and constraints are detailed in natural language. Although the study reports failure rates stemming from a range of issues, the authors estimate that 25% of failures occur since the agent "fails to understand the domain-specific knowledge or rules". A recent work by Li et al. (2025) offers a framework for evaluation of policy adherence in multi-step single-turn tasks in five diverse domains. The authors report relatively low pass-rate across agents, varying from 31% to 69%. Both studies introduce policies to the agent in its prompt, instructing the LLM to follow the guidelines while satisfying user requirements.
Contrary to the pro-active policy enforcement above, another line of studies take the post-hoc reflection approach, by designing guardrails that check if any action taken by an agent violates a predefined policy (Huang et al., 2024; Rothkopf et al., 2024; Hua et al., 2024). While this strategy is often easier for implementation (e.g., detecting that an unauthorized refund was issued is simpler than preventing it), it may fall short of the strict demands of enterprise-grade business flow automation, particularly in mission-critical contexts.
Additional studies deal with related, yet not strictly relevant fields of agent safety, security and robustness to adversarial attacks (Levy et al., 2024; Nakash et al., 2024, 2025; Chennabasappa et al., 2025). These works evaluate LLM agents’ resilience to malicious requests (such as exposing admin credentials) and propose mitigation strategies. Another thread of work explores the benefits of model alignment for meeting domain-specific regulations by tuning a model against request-response pairs that comply with company guidelines (Achintalwar et al., 2024). However, this approach better suits the scenario of QA agents, contrary to business process automation, where actions ultimately mutate the system state (e.g., book a flight), and agent is required to perform complex reasoning, based on the state and conversation history.
Our work is the first, to the best of our knowledge, to propose an automatic pipeline for pro-active policy enforcement in agentic flows, involving only minimal human intervention.
3 Methodology
We present an end-to-end pipeline for pro-active enforcement of policy adherence in a realistic setting, assuming that policies are provided as free-form natural language documents, and that function specifications (the agent’s toolkit) and the system data schema are available. The offline buildtime process consists of two main components, as illustrated in Figure 1, which we describe in detail below.
3.1 Tool-Policy Mapper
The primary goal of the mapper is to transform an (often) lengthy and noisy natural language policy document into a compact, structured representation by linking clearly formulated policy statements to relevant tools from the provided toolkit. Business process guidelines are typically authored by domain or legal experts, and mapping them to tools requires advanced reading comprehension skills. As a concrete example, consider the following policy from the $\tau$-bench Airlines domain (verbatim from Yao et al. (2024), enumerated for clarity):
(1) Flight reservations can be canceled within 24 hours of booking, or if the airline canceled the flight. Otherwise, (2) basic economy or economy flights can be canceled only if travel insurance is bought […], and (3) explicit customer confirmation should be obtained prior to canceling a flight."
All three atomic policies above should be mapped to the cancel_reservation() function. The tool’s invocation is then conditioned on cancellation eligibility, as verified by the corresponding guard. In addition to recording each policy’s name and description, the mapper also extracts supporting textual spans (references) from the source document. It then generates multiple example requests for policy compliance and violation, a step shown to be effective for both human evaluation of the mapper’s output and for facilitating efficient guard generation (see Section 3.2 for details). The mapper is also guided to split policies associated with a given tool into smaller, logically distinct units – similarly to how a human would break down a lengthy policy description into clear, fine-grained guidelines. For instance, policies concerning flight cancellations can be divided into: (a) rules about cancellation time windows for full or partial refunds, (b) policies specific to canceling individual segments within connecting flights, etc. Table 1 shows an example of (partial) mapper outcome for a policy item, related to the cancel_reservation() tool.
policy Flexible Cancellation Policy